Here is the Agri-Data, mention the name of dependent and independent variables.


Part A: Prepare the line of regression Y on X.


Part B: Calculate r2 (coefficient of determination). Interpreted the byx.


District_2 Rapeseed & Rapeseed &

Mustard Production Mustard Area (Hectare)

in Metric tonnes

UTTAR KASHI 437 87

CHAMOLI 193 407

RUDRA PRAYA 154 325

TEHRI GARHW 466 848

DEHRADUN 139 226


1
Expert's answer
2021-09-01T15:55:07-0400

Data




Rapeseed & Mustard production in Metric Tones is dependent variable = Y

Rapeseed mustard area production is an independent variable = X

Part A

By using Excel




"y=a + bx"

=INTERCEPT(A2:A6,B2:B6)

"a = -2.28"

=SLOPE(A2:A6,B2:B6)

"b= 0.52\\

y=0.52-2.28x"

Part B

=RSQ(A2:A6,B2:B6)

"R^2=0.981"

The relation between Y and X is very strong. 98% of the variance of the dependent variable being studied is explained by the variance of the independent variable.

The graphical method with trend line.